OpenStudy (anonymous):

write an equation. the length of a rectangle is 4 inches more than it's width. the perimeter is 96.

OpenStudy (anonymous):

p = 2(w+l) l = w+4, l=x+4, w=x p = 2(x+4+x) p=2(2x+4) p=(4x+8) 96=4x+8 4x=96-8 4x=88 x=22 w=22, l=26 96=2(22+26) 96=2(48) 96=96 qed